OpenStudy (anonymous):

N is a natural no formed by writing ,in the ascending order,the first 1002 whole numbers one after another.Find the remainder when N is divided by 9

OpenStudy (anonymous):

Is the remainder 2??

OpenStudy (anonymous):

Remainder of (N/9) = Rem(Sum of the digits of N/9) ... U CAN CHECK IT. Now N=12345......999 1000 10001 UPTO 9999 no rema will be 0 for 1000 rem= sum of digit=1 for 1001 remainder= 2 therefoe ans= 1+2 =3

OpenStudy (anonymous):

He gr8... thanks a lot vicky.

Parth (parthkohli):

We can calculate the sum of digits from 0 to 1001. \( \color{Black}{\Rightarrow \Large {1002 \over 2}({0 + 1001 )}}\) Arithmetic sequence ^

Parth (parthkohli):

\( \color{Black}{\Rightarrow 501(1001) = ......1 }\)

OpenStudy (anonymous):

Hey lil correction: my ans is for digits upto 1001 for ur ques ... 1002 rem (when div by 9)= 3 net ans= 1+2+3=6

Parth (parthkohli):

501501 ===> 5 + 1 + 1 + 5 = 1 + 2 = 3 3 is not divisible by 9, but 9 - 3 = 6

Parth (parthkohli):

@Aprajita111 Did you understand?

OpenStudy (anonymous):

Yes

Parth (parthkohli):

Great!